rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_red
icon_orange
icon_red
icon_red
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#4177 — FS#8111 — Bordeaux optical loop

Attached to Project— ADSL / SDSL
Maintenance
Backend / Core
CLOSED
100%
Works are processing on a fiber.
The security has taken over.

There's no service cutoff.


Date:  Sunday, 03 March 2013, 13:34PM
Reason for closing:  Done
Comment by OVH - Saturday, 23 February 2013, 16:48PM

We intervened on the site of Talence in order to replace an optical garter . In fact the garter was giving a non conform optical level. After this replacement everything was in order again.

The securisation is operational again.